Evaluate the following expressions.
Note: Your answer must be in EXACT form: it cannot contain decimals. It must be either an integer or a fraction. If the answer involves a square root write it as sqrt . For instance, the square root of 2 should be written as sqrt(2).
sin(-2pi/3) = ?
cos(pi/2) = ?
tan(pi/6) = ?
cot(pi/2) = ?
sec(4pi/3) = ?
csc(-pi/6) = ?

Answer by stanbon(48558)

Evaluate the following expressions.
Note: Your answer must be in EXACT form: it cannot contain decimals. It must be either an integer or a fraction. If the answer involves a square root write it as sqrt . For instance, the square root of 2 should be written as sqrt(2).
sin(-2pi/3) = -sin(2pi/3) = -sin(120 degrees) = -(1/2)
cos(pi/2) = 0
tan(pi/6) = 1/sqrt(3) or sqrt(3)/3
cot(pi/2) = cos(pi/2)/sin(pi/2) = 0/1 = 0
sec(4pi/3) = 1/cos(4pi/3) = 1/cos(240)= 1/-cos(60) = -2
csc(-pi/6) = -csc(pi/6) = -1/sin(pi/6) = -2
===============================================================
Cheers,
Stan H.